# Probing Reheating Phase via Non-Helical Magnetogenesis and Secondary Gravitational Waves

### v1

@@ II.2.2 Evolution During reheating:
Many existing cosmological magnetogenesis models assume that, following inflation, fields on super-horizon scales experience a radiation-like redshift, with the magnetic energy density scaling as $\rho_{\rm em}\propto a^{-4}$ Campanelli (2009); Jain et al. (2014); Adshead et al. (2016); Kahniashvili et al. (2017); Tripathy et al. (2022); Li and Zhang (2022). However, it is important to note that the individual evolution of the electric and magnetic field components, $E^{2}$ and $B^{2}$, may deviate from this assumption, as their behavior depends sensitively on the conductivity of the universe during reheating Benevides et al. (2018); Kobayashi and Sloth (2019); Haque et al. (2021b). This nontrivial evolution can be attributed to Faraday induction, through which electric energy can be converted into magnetic energy, and vice versa, depending on the field strengths Haque et al. (2021b); Kobayashi and Sloth (2019).
In non-helical magnetogenesis scenarios, it is typically observed that either the magnetic or electric field strength dominates at the end of inflation, as pointed out before about the Eqs. (14) and (15). Due to this hierarchy between the electric and magnetic energy densities, the subsequent evolution is strongly influenced by the conductivity. In cases where the electric field dominates over the magnetic field at the end of inflation (for $n<-1/2$), the electric energy is gradually converted into magnetic energy, thereby amplifying the magnetic field strength Haque et al. (2021b). Conversely, when the magnetic field dominates (for $n>1/2$), energy is transferred from the magnetic field to the electric field.
In the high-conductivity limit, rapid charge separation generates counter-electric fields that effectively cancel the electric component, leaving only the magnetic field. In this case, the magnetic field evolves adiabatically, and its energy density redshifts as $\rho_{\rm B}\propto a^{-4}$. Although our primary interest lies in modes that are far outside the horizon at the end of reheating, the role of conductivity for such superhorizon modes remains subtle. Conductivity arises in a causal process; however, for the super-horizon modes, causality prevents different regions from communicating, raising the possibility that these large-scale modes might not “feel” the effects of high conductivity, even though the universe as a whole is highly conductive. If these modes indeed remain unaffected by conductivity, the dynamics involving energy conversion between the electric and magnetic fields would still apply during reheating. But for this paper, we simply consider two limiting scenarios in our analysis: one with zero electrical conductivity and another with infinite electrical conductivity during the reheating phase.
The case of negligible conductivity during reheating
If we assume zero electrical conductivity in the universe during the reheating period, both electric and magnetic fields are present, with the strength of the fields depending on their coupling to the background. As shown in Eqs.(16) and (17), for $n>1/2$, we obtain a stronger magnetic field, and for $n<-1/2$, the electric field’s strength is dominant over the magnetic field at early stages. We consider both scenarios, and their evolution during reheating is governed by the following.
$\mathcal{P}_{B}>>\mathcal{P}_{E}$: The modes of interest in this context have a scale of approximately 1 $1\mbox{ Mpc}^{-1}$, which is well outside the Hubble radius at the end of reheating. Therefore, we can consider the limit where $k(\eta-\eta_{end})<<1$, and in this limit, the equations (22a) simplify to:
$a_{\mathrm{re}}^{4}\mathcal{P}_{B}(k,\eta_{\rm re})\simeq\mathcal{B}-(\mathcal{B}-\mathcal{E})k^{2}\eta_{\rm re}^{2}$ (26)
$a_{\mathrm{re}}^{4}\mathcal{P}_{E}(k,\eta_{\rm re})\simeq\mathcal{E}+(\mathcal{B}-\mathcal{E})k^{2}\eta_{\rm re}^{2}$ (27)
It is worth noting that for positive values of $n$, the magnetic field is stronger than the electric field, as indicated in Eq. (16). Furthermore, from Eqs. (24) and (25), it is evident that $\mathcal{B}\gg\mathcal{E}$. As demonstrated above, the efficiency of energy transfer from magnetic to electric components is notably sensitive to the duration of the reheating phase. We will explore this dependence in greater detail in a subsequent section.
$\mathcal{P}_{E}>>\mathcal{P}_{B}$: In a similar manner to what was stated previously, if we take into account the negative value of $n$, we observe that the electric field surpasses the magnetic field strength at the end of inflation. Consequently, during reheating on the superhorizon scale limit, the spectrum evolves into,
$a_{\mathrm{re}}^{4}\mathcal{P}_{E}(k,\eta_{\rm re})\simeq\mathcal{B}-(\mathcal{B}-\mathcal{E})k^{2}\eta_{\rm re}^{2}$
$a_{\mathrm{re}}^{4}\mathcal{P}_{B}(k,\eta_{\rm re})\simeq\mathcal{E}+(\mathcal{B}-\mathcal{E})k^{2}\eta_{\rm re}^{2}.$ (28)
In this instance, we note that the electric field energy density is converted into magnetic field energy density. We have also found that the conversion of the energy density is proportional to the energy difference of the two fields, and it is directly related to the total duration of the reheating periods. It means a longer duration of the reheating periods allows more conversion of the energy density. Effectively, at the super-horizon scale, the strength of the magnetic field also increases significantly. Such a conversion mechanism turns out to be helpful for the $n<-1/2$ case, when the magnetic field strength at large scale is low at the end of inflation, as it yields a strong magnetic field at present without any strong coupling as well as backreaction issues.
Based on the preceding discussion, it is evident that if we consider zero (or small) conductivity during the reheating phase, the individual components of the electromagnetic field, namely $B^{2}$ and $E^{2}$, undergo non-trivial evolution. Moreover, the field strength is directly linked to the reheating parameters, specifically the reheating temperature $T_{\rm re}$ and the average equation of state $w_{re}$. If we wish to express how the comoving magnetic field evolves as a function of these reheating parameters, we must carefully account for the energy transfer dynamics between the electric and magnetic components, and we have the following
$a_{\mathrm{re}}^{4}\mathcal{P}_{B}(k,\eta_{re})=\left\{\begin{matrix}\mathcal{B}-(\mathcal{B}-\mathcal{E})\left(\frac{k}{k_{\rm e}}\right)^{2}\mbox{Exp}\left[N_{\mathrm{re}}(1+3w_{\rm re})\right]&\mbox{for $n>\frac{1}{2}$}\\ \mathcal{E}+(\mathcal{B}-\mathcal{E})\left(\frac{k}{k_{\rm e}}\right)^{2}\mbox{Exp}\left[N_{\mathrm{re}}(1+3w_{\rm re})\right]&\mbox{ for $n<-\frac{1}{2}$}\end{matrix}\right.$ (29)
Here $N_{\mathrm{re}}$ is the e-folding number of the reheating phase, which can be directly related to the reheating temperature through Eq.(8) Martin and Ringeval (2010); Maity and Saha (2018); Haque and Maity (2023b).
Figure 1: The figures mentioned above illustrate the present-day magnetic field strength at the 1 Mpc length scale, denoted as $B_{0}$ (in Gauss), as a function of the equation of state $w_{re}$. Different colors represent various reheating temperatures $T_{\rm re}$ (in GeV). The right plot is $n=-2$, resulting in a pronounced blue-tilted magnetic spectrum ($n_{\mathrm{B}}=2$). In contrast, the left figure represents the case of $n=2$, yielding a scale-invariant magnetic spectrum with $n_{\mathrm{B}}=0$. In both figures, the cyan-shaded region denotes an excluded area due to the upper limit on the present-day magnetic field strength at the 1 Mpc scale, as derived from constraints imposed by the Cosmic Microwave Background (CMB) Paoletti et al. (2022b); Zucca et al. (2017).
The case of infinite conductivity during reheating:
If we assume an infinite conductivity limit during reheating, the only remaining field is the magnetic field, and its energy density evolution follows the relation $\rho_{em}=\rho_{B}\propto a^{-4}$. The energy density dilutes due to the expansion of the universe, and the comoving magnetic power spectrum freezes after inflation ends as any conversion or production ceases to exist during this period. Thus, at the superhorizon scale, after the end of reheating, the magnetic power spectrum is defined as
$\begin{matrix}a_{\mathrm{re}}^{4}\mathcal{P}_{B}(k,\eta_{\rm re})&=\frac{k^{4}}{8\pi}\frac{2^{2n+1}\Gamma^{2}\left(n+\frac{1}{2}\right)}{\pi^{2}}\left(\frac{k}{k_{f}}\right)^{-2n}\hskip 28.45274pt\mbox{ for $n>\frac{1}{2}$}\\ a_{\mathrm{re}}^{4}\mathcal{P}_{B}(k,\eta_{\rm re})&=\frac{k^{4}}{8\pi}\frac{2^{2n-1}\Gamma^{2}\left(n-\frac{1}{2}\right)}{\pi^{2}}\left(\frac{k}{k_{f}}\right)^{2-2n}\ \hskip 28.45274pt\mbox{ for $n<-\frac{1}{2}$}\end{matrix}$ (30)
Constraining Parameters from CMB Temperature Fluctuations ($\delta T/T\simeq 10^{-5}$):
The evolution of the universe after inflation is critical for determining the present-day magnetic field strength generated during this era. During reheating, the background energy density (inflaton energy density $\rho_{\phi}$) evolves as $a^{-3(1+w_{\rm re})}$. In contrast, the electromagnetic (EM) energy density $\rho_{\mathrm{em}}$ evolves as $a^{-4}$ due to background expansion. Thus, the EM energy density fluctuation $\delta_{\mathrm{EM}}=\rho_{\mathrm{EM}}/\rho_{\phi}$ scales as $\delta_{\mathrm{EM}}\propto a^{(3w_{\rm re}-1)}$, making it sensitive to both the fluid’s nature (defined by $w_{\rm re}$) and the reheating duration (set by the reheating temperature $T_{\rm re}$).
To align with CMB observations, the temperature fluctuations should satisfy $\delta T/T\leq 10^{-5}$. Assuming all temperature fluctuations arise from the EM energy density (although this is a very crude approximation), the EM energy density at the end of reheating must adhere to this constraint. Due to high conductivity following reheating, any electric field dissipates quickly, leaving only the magnetic component relevant when $n>1/2$ (where $n$ is the coupling constant). This analysis excludes MHD effects, which are significant only for sub-horizon modes and may have changed the bounds.
Based on these considerations, we derive the constraint:
$\frac{2^{2n+1}\Gamma^{2}(n+1/2)}{24\pi^{3}(4-2n)}\left(\frac{H_{{}_{\mathrm{I}}}}{M_{\rm P}}\right)^{2}\left\{1-\left(\frac{k_{*}}{k_{\rm e}}\right)^{4-2n}\right\}\left(\frac{a_{\mathrm{e}}}{a_{\mathrm{re}}}\right)^{1-3w_{\rm re}}\leq 4\frac{\delta T}{T},$ (31)
where this bound is effective only for $w_{re}>1/3$ due the factor $(a_{\mathrm{e}}/a_{\mathrm{re}})^{1-3w_{re}}$. From this, we constrain the reheating dynamics and determine the maximum possible present-day magnetic field strength, denoted $B_{0}$.
EoS $w_{re}=0.5$ $w_{re}=0.6$
$n_{\mathrm{B}}$ 0.01 0.2 0.4 0.6 0.01 0.2 0.4 0.6
$T_{\rm re}$ (GeV) $9.09\times 10^{5}$ $1.84\times 10^{3}$ $215.5$ $48.9$ $1.68\times 10^{9}$ $2.7\times 10^{7}$ $6.52\times 10^{6}$ $2.2\times 10^{6}$
$B_{0}$ (G) $1.47\times 10^{-9}$ $1.4\times 10^{-11}$ $5.0\times 10^{-14}$ $1.51\times 10^{-16}$ $1.475\times 10^{-9}$ $1.407\times 10^{-11}$ $5.04\times 10^{-14}$ $1.56\times 10^{-16}$
Table 1: Upper bounds on reheating temperature consistent with CMB temperature fluctuations (Eq. (31)) and the corresponding present-day magnetic field strength $B_{0}$ in Gauss.
In Table (1), we present the lower bounds on the reheating temperature for $w_{re}=0.5$ and $w_{re}=0.6$, across various values of the magnetic spectral index $n_{\mathrm{B}}=4-2n$. We also provide the corresponding present-day magnetic field strength in Gauss, defined at a scale of $1~{}\text{Mpc}^{-1}$. These bounds are derived under certain assumptions discussed earlier.
@@
In particular, we neglect the magnetohydrodynamic (MHD) evolution of sub-horizon modes, which are most significantly affected by MHD processes. Therefore, for blue-tilted magnetic spectra, the results presented here are expected to be substantially modified in a more realistic scenario. However, since a full treatment of MHD dynamics is beyond the scope of this work, we present these bounds as indicative estimates.
[…]
In Fig. (1), we depict the strength of the present-day magnetic field $B_{0}$ as a function of the equation of state $w_{re}$, with $w_{re}$ ranging from 0 to 1. Four different colors represent four distinct reheating temperatures. In the left plot, we assume a specific coupling $n=2$, which yields a scale-invariant magnetic spectrum, while the electric spectrum scales as $\mathcal{P}_{\rm E}\propto k^{2}$. Depending on the duration of reheating, a significant conversion of energy occurs from the magnetic field to the electric field ($\rho_{\rm B}\to\rho_{\rm E}$). However, for this particular coupling ($n=2$), the comoving magnetic energy density remains conserved on large scales.
We observe that as the equation of state increases, the present-day magnetic field strength also gradually increases. This is due to the rate of dilution: when $w_{re}>1/3$, the background energy density dilutes faster than the radiation energy density, and the universe spends less time in reheating compared to scenarios with $w_{re}<1/3$. Hence, the relative dilution of the magnetic field compared to the background is diminished, causing the effective present-day magnetic field strength to be larger for a higher equation of state.
For a scale-invariant magnetic spectrum ($n_{\mathrm{B}}=0$), the maximum allowed equation of state (EoS) parameter during reheating is approximately $w_{re}\simeq 0.5$ across a wide range of reheating temperatures, as illustrated in Fig. 1. Higher values of $w_{re}$ are disfavoured, as they lead to magnetic field strengths at the $1~{}\text{Mpc}^{-1}$ scale that exceed the current upper limits set by CMB observations. In contrast, for a blue-tilted magnetic spectrum, this constraint is significantly relaxed, as shown in the right panel of Fig. 1.
In the right panel, we plot $B_{0}$ against $w_{re}$ for different reheating temperatures $T_{\rm re}$, considering a different coupling relationship ($n=-2$), which results in a scale-dependent magnetic field. In this case, the magnetic power spectrum follows a $\mathcal{P}_{\rm B}(k,\eta_{0})\propto k^{2}$ behavior, leading to a stronger electric field compared to the magnetic field. In the absence of electrical conductivity, we observe significant energy conversion from the electric field to the magnetic field, with the extent of conversion depending on the duration of reheating.
Lowering the reheating temperature increases the duration of the reheating phase, allowing greater energy transfer and leading to higher magnetic field strengths. Interestingly, we find that inflationary production alone, without prolonged reheating, is insufficient to generate magnetic fields strong enough to match present-day observations for coupling types with $n<-1/2$.
However, when prolonged reheating with negligible electrical conductivity is considered, efficient energy transfer from the electric to the magnetic field becomes possible, thereby enhancing the magnetic field strength to observationally viable levels. For example, with $w_{re}=0.5$ and $T_{\rm re}\sim 10^{-2}~{}\mathrm{GeV}$, even for couplings with $n<-1/2$, the magnetic field strength at $1~{}\mathrm{Mpc}$ can reach up to $B_{0}\sim 10^{-14}~{}\mathrm{G}$, which lies well within the current observational bounds.
In contrast, if we impose the CMB constraints on temperature fluctuations discussed above, then for $w_{re}=0.5$, the lowest allowed reheating temperature is $T_{\rm re}\sim 9.1\times 10^{5}~{}\mathrm{GeV}$. In this case, the corresponding present-day magnetic field strength is $B_{0}\sim 6.4\times 10^{-23}~{}\mathrm{G}$, which is about $10^{12}$ times larger than the scenarios where we have no such conversion of the energy from electric field to magnetic field. In radiation-like reheating scenarios with negligible electrical conductivity, for a reheating temperature of $T_{\rm re}=10^{-2}~{}\mathrm{GeV}$, the maximum possible conversion of electric field energy into magnetic field energy is allowed. In this case, we find that the present-day magnetic field strength can reach up to $B_{0}\sim 2.6\times 10^{-18}~{}\mathrm{G}$.
On the other hand, for matter-like reheating scenarios ($w_{re}=0$) with the same reheating temperature $T_{\rm re}=10^{-2}~{}\mathrm{GeV}$, even after significant conversion from electric to magnetic fields, the present-day magnetic field strength is much lower, approximately $B_{0}\sim 3.7\times 10^{-30}~{}\mathrm{G}$. Although a matter-like reheating phase leads to a longer reheating duration, allowing more efficient conversion of electric energy into magnetic energy, the extended post-inflationary expansion results in greater dilution of the magnetic energy density, causing the present-day magnetic field strength to be significantly lower despite the effective energy transfer.
@@
Thus, the post-inflationary evolution of the gauge field plays a crucial role in determining the present-day magnetic field strength. Furthermore, the electrical conductivity of the Universe during reheating significantly influences the final magnetic field amplitude.

### v2

@@ II.3 Evolution During reheating phase:
    During the post-inflationary phase, if there is no further gauge field production, the electromagnetic energy density scales as $\rho_{\rm em}\propto a^{-4}$ Campanelli (2009); Jain et al. (2014); Adshead et al. (2016); Kahniashvili et al. (2017); Tripathy et al. (2022); Li and Zhang (2022). However, it is important to note that the individual evolution of the electric and magnetic field components, $E^{2}$ and $B^{2}$, may deviate from each other, as their behavior depends sensitively on the electrical conductivity of the universe during reheating Benevides et al. (2018); Kobayashi and Sloth (2019); Haque et al. (2021b). This nontrivial evolution can be attributed to Faraday induction, through which electric energy can be converted into magnetic energy, and vice versa, depending on the field strengths Haque et al. (2021b); Kobayashi and Sloth (2019). For example, the case we just discussed in the above Eq.24a, where the Faraday effect is evident. However, it would be pronounced in large-scale limit $k\eta\rightarrow 0$, when the above equations are approximated as,
    $\mathcal{P}_{B}=\frac{1}{a^{4}}(\mathcal{B}+(\mathcal{B}-\mathcal{E})k^{2}(\eta-\eta_{\rm e})^{2})\simeq\frac{\mathcal{B}}{a^{4}}+\frac{\mathcal{B}-\mathcal{E}}{H^{2}a^{6}}k^{2},$ (27)
    $\mathcal{P}_{E}=\frac{1}{a^{4}}(\mathcal{E}-({\mathcal{B}-\mathcal{E}})k^{2}(\eta-\eta_{\rm e})^{2})\simeq\frac{\mathcal{E}}{a^{4}}-\frac{\mathcal{B}-\mathcal{E}}{H^{2}a^{6}}k^{2},$ (28)
    Where, we utilize the relation $\eta-\eta_{\rm e}\sim 1/(aH)$. The above equation clearly indicates how the Faraday effect transfers energy from electric to magnetic modes and vice versa with time behavior $\mathcal{P}_{B}\propto 1/a^{6}H^{2}$, which has been discussed earlier in the literature Kobayashi (2014); Haque et al. (2021b).
    In what follows, we focus on the intermediate regime 2.2<n<0-2.2<n<0, corresponding to $0.4<n_{\mathrm{E}}<4$, where both strong coupling and backreaction problems are absent. In this case, as stated earlier electric field is stronger than the magnetic field at the end of inflation, $\mathcal{E}\gg\mathcal{B}$. The Faraday effect, therefore, leads to an additional enhancement of the magnetic field, and at the end of reheating, one obtains,
    $\mathcal{P}_{E}(k,\eta_{\rm re})\simeq\frac{1}{a_{\mathrm{re}}^{4}}(\mathcal{E}+(\mathcal{B}-\mathcal{E})k^{2}\eta_{\rm re}^{2})\simeq\frac{\mathcal{E}}{a_{\mathrm{re}}^{4}}$ (29)
    $\mathcal{P}_{B}(k,\eta_{\rm re})\simeq\mathcal{B}-(\mathcal{B}-\mathcal{E})k^{2}\eta_{\rm re}^{2}\simeq\frac{\mathcal{E}k^{2}}{a_{\mathrm{re}}^{6}H_{\rm re}^{2}}.$ (30)
    After the end of reheating, assuming the universe is dominated by the relativistic plasma with infinite conductivity, the electric field quickly decays, and the magnetic field evolves adiabatically till the present day. The present-day magnetic field $B_{0}$ can be calculated as
    $a_{0}^{4}\mathcal{P}^{0}_{B}(k,\eta_{\rm re})\simeq\frac{\mathcal{E}k^{2}}{a_{\mathrm{re}}^{2}H_{\rm re}^{2}}\implies B_{0}\simeq\frac{\sqrt{\mathcal{E}}k}{a_{\mathrm{re}}H_{\rm re}{a_{0}^{2}}}$ (31)
    Here, we have observed that due to the Faraday induction effect, there is significant conversion of electric field energy into magnetic field energy. To highlight this effect, we present some numerical comparisons for two extreme cases of electrical conductivity. For the zero electrical conductivity (EC) case with coupling parameter $n=-2.0\,(n_{\mathrm{E}}=0)$, where the magnetic field energy density scales as $k^{2}$, we find that for a reheating temperature $T_{\rm re}=10^{-2}\,\mathrm{GeV}$ and equation of state $w_{re}=0$, the present-day magnetic field strength is $B_{0}(1\,\mathrm{Mpc}^{-1})\simeq 1.86\times 10^{-27}\,\mathrm{G}$. In contrast, for the same parameters but assuming infinite EC, the strength drops to $B_{0}(1\,\mathrm{Mpc}^{-1})\simeq 8.85\times 10^{-40}\,\mathrm{G}$. Similarly, for the same reheating temperature with $w_{re}=1/3$, we find that the present-day magnetic field strength in the zero EC case is $B_{0}(1\,\mathrm{Mpc}^{-1})\simeq 1.27\times 10^{-15}\,\mathrm{G}$, while the infinite EC case yields $B_{0}(1\,\mathrm{Mpc}^{-1})\simeq 7.3\times 10^{-34}\,\mathrm{G}$. Therefore, it is evident that for a fixed reheating temperature, the zero conductivity scenario leads to a significant enhancement in the magnetic field strength, up to a factor of $10^{14}-10^{19}$ compared to the infinite conductivity case, due to efficient electric-to-magnetic energy conversion.
Constraining the Reheating parameter from the Backreaction at the end of Reheating:
Before we embark on the production of gravitational waves, let us derive a generic constraints on the reheating parameter space due to additional production of gauge field. As we now consider a general reheating history, it is important to account for the possibility that the equation of state (EoS) parameter during reheating satisfies $w_{re}>1/3$. In such cases, the background energy density dilutes faster than the electromagnetic (EM) energy density. Specifically, the inflationary energy density evolves as $\rho_{\phi}(\eta)\propto a^{-3(1+w_{re})}$, whereas the electromagnetic energy density evolves as $\rho_{\rm em}(\eta)\propto a^{-4}$. This implies that the fractional energy density of the EM field increases during reheating according to $\delta\rho_{\rm em}(\eta)\equiv\frac{\rho_{\rm em}(\eta)}{\rho_{\phi}(\eta)}\propto a^{3w_{re}-1}\,.$ Therefore, depending on the reheating dynamics, there may arise a situation in which the EM energy density becomes comparable to or even exceeds the background energy density.
This scenario is particularly constrained by Big Bang Nucleosynthesis (BBN). If the universe is dominated by magnetic field energy at the onset of BBN, it might disrupt the formation of light elements. Since the predictions of BBN are in excellent agreement with observations, we must impose a consistency condition that the total EM energy density at the end of reheating remains subdominant to the background radiation energy density, $\rho_{\rm em}(\eta_{\rm reh})<\rho_{r}(\eta_{\rm reh})\,$. Imposing this condition leads to nontrivial constraints on the reheating dynamics.
These constraints are further modified depending on the post-inflationary conductivity of the universe. Let us first consider the case of high electrical conductivity. In this limit, only the magnetic field survives during reheating. As discussed previously, and within the allowed parameter range 2.2<n<0-2.2<n<0, the magnetic power spectrum is blue-tilted and evolves adiabatically, as described by Eq. (19). In this high-conductivity regime, the fractional energy density of the magnetic field evolves as $\delta\rho_{B}(\eta)\equiv\frac{\rho_{B}(\eta)}{\rho_{\phi}(\eta)}\propto a^{3w_{re}-1}\,$, similar to the general EM case. Thus, for sufficiently large $w_{re}$, the magnetic field can become dynamically significant and potentially dominate the energy budget, leading to tight constraints on the reheating duration and EoS.
$\delta\rho_{B}(\eta_{\rm re})\simeq\left(\frac{H_{{}_{\mathrm{I}}}}{M_{\rm P}}\right)^{2}\frac{2^{3-n_{\mathrm{E}}}\Gamma^{2}\left(\frac{3-n_{\mathrm{E}}}{2}\right)(1-(k_{*}/k_{\rm e})^{n_{\mathrm{E}}+2})}{24\pi^{3}(n_{\mathrm{E}}+2)}\exp[N_{\mathrm{re}}(3w_{re}-1)]$ (32)
Figure 1: In this figure, we present the lowest allowed reheating temperature (in GeV) as a function of the reheating equation of state parameter $w_{re}$, over a wide range. The three different colors correspond to three distinct values of the electric spectral index: $n_{\mathrm{E}}=0.01$ (blue), $n_{\mathrm{E}}=0.1$ (red), and $n_{\mathrm{E}}=0.2$ (magenta). Solid lines represent the scenario with zero electrical conductivity (EC), while dashed lines correspond to the high EC limit. Throughout the analysis, we fix the inflationary energy scale at $H_{\rm inf}\simeq 10^{-5}M_{\rm Pl}$.
On the other hand, in the zero electrical conductivity limit, both electric and magnetic fields are present during the reheating era. We assume that the electric field decays only after reheating completes, when the universe becomes fully dominated by radiation. Therefore, before the end of reheating, both fields contribute to the total electromagnetic energy density. To consistently avoid backreaction in this regime, it is essential to account for the combined influence of both the electric and magnetic fields. Consequently, in the zero conductivity limit, the electromagnetic fractional energy density must be defined by including contributions from both components.
$\delta\rho_{em}(\eta_{\rm re})\simeq\left(\frac{H_{{}_{\mathrm{I}}}}{M_{\rm P}}\right)^{2}\frac{2^{5-n_{\mathrm{E}}}}{24\pi^{3}}\left[\frac{\Gamma^{2}\left(\frac{5-n_{\mathrm{E}}}{2}\right)(1-(k_{*}/k_{\rm e})^{n_{\mathrm{E}}})}{n_{\mathrm{E}}}+\frac{\Gamma^{2}\left(\frac{3-n_{\mathrm{E}}}{2}\right)(1-(k_{*}/k_{\rm e})^{n_{\mathrm{E}}+2})}{4(n_{\mathrm{E}}+2)}\right]\exp[N_{\mathrm{re}}(3w_{re}-1)]$ (33)
As seen from Eqs.(32) and (33), the fractional electromagnetic energy density increases with time only in reheating scenarios with $w_{re}>1/3$. For example, in the high-conductivity limit with $w_{re}=0.5$, the minimum reheating temperature required to prevent the magnetic field from dominating before the onset of the radiation-dominated era is $T_{\rm re}^{\rm min}\sim 10^{-2}\,\mathrm{GeV}$. In contrast, for the same equation of state but in the zero electrical conductivity limit, the corresponding minimum reheating temperature is significantly higher, $T_{\rm re}^{\rm min}\sim 3\times 10^{2}\,\mathrm{GeV}$.
In Fig. 1, we present the lowest allowed reheating temperature, denoted by $T_{\rm re}^{\rm min}$ (in GeV), as a function of the reheating equation of state parameter $w_{re}$, for three different values of the electric spectral index: $n_{\mathrm{E}}=0.01$, $0.1$, and $0.2$. These parameter choices ensure that the model remains free from both backreaction and strong coupling problems during inflation as well as during reheating. As previously discussed, for $w_{re}>1/3$, the background energy density dilutes faster than the electromagnetic (EM) energy density, allowing the latter to potentially dominate at later stages of reheating. The curves in the figure indicate the threshold values of $T_{\rm re}$ below which the EM field would dominate over the background energy density before reheating completes.
In this plot, solid lines represent the zero electrical conductivity limit, while dashed lines correspond to the high conductivity limit. As observed, the constraints in the high conductivity scenario are less sensitive to the spectral index $n_{\mathrm{E}}$, since only the magnetic field contributes, and the total magnetic energy density, being strongly blue-tilted, is relatively insensitive to variations in $n_{\mathrm{E}}$. In contrast, in the zero conductivity case, the majority of the EM energy budget is due to the electric field, whose energy density is highly sensitive to $n_{\mathrm{E}}$ for the chosen parameter range.
Figure 2: In this figure, we show the present-day magnetic field strength evaluated at the scale $1\,\mathrm{Mpc}^{-1}$ as a function of the reheating equation of state parameter $w_{re}$. The three different colors represent three distinct values of the reheating temperature. Solid lines correspond to $n_{\mathrm{E}}=0.01$, while dashed lines indicate $n_{\mathrm{E}}=0.20$. The left panel illustrates the case of zero electrical conductivity during reheating, whereas the right panel corresponds to the infinite conductivity limit. In both cases, we impose the condition that the total electromagnetic (EM) energy density remains subdominant to the background energy density throughout reheating, thereby ensuring that the universe does not enter a magnetic field-dominated phase before the onset of Big Bang Nucleosynthesis (BBN).
@@
In this work, we neglect the magnetohydrodynamic (MHD) evolution of sub-horizon modes, which are most susceptible to MHD effects. Consequently, for blue-tilted magnetic spectra, the estimates provided here are expected to be significantly altered in more realistic scenarios. However, since a full treatment of MHD dynamics lies beyond the scope of this study, we present our results as indicative bounds.
Using these bounds, we analyze the behavior of the present-day magnetic field under two limiting cases, zero electrical conductivity and infinite electrical conductivity during reheating. In Fig. 2, we plot the present-day magnetic field strength as a function of the reheating equation of state $w_{re}$, for three different reheating temperatures: $T_{\rm re}=1\,\mathrm{GeV},\,10^{5}\,\mathrm{GeV},\,\text{and }10^{8}\,\mathrm{GeV}$.
In the left panel of Fig. 2, corresponding to the zero conductivity limit, we observe that a significant portion of the electric field energy is converted into magnetic field energy during reheating. As a result, the present-day magnetic field amplitude is considerably enhanced. In contrast, the right panel corresponds to the infinite conductivity limit, where no such conversion takes place. Consequently, the magnetic field strength today is much smaller compared to the zero conductivity case.
We further observe that increasing the equation of state $w_{re}$ and decreasing the reheating temperature $T_{\rm re}$ both tend to enhance the present-day magnetic field strength. This enhancement is driven by the dilution of magnetic energy due to cosmic expansion. For larger $w_{re}$, the background energy density redshifts more rapidly, allowing the Universe to reach lower reheating temperatures over a shorter timescale. As a result, magnetic fields experience less dilution from expansion, leading to a stronger residual amplitude today.
Moreover, we find that for more blue-tilted electric spectra, the magnetic field strength at large scales decreases. This is because the magnetic spectrum becomes increasingly blue-tilted with growing $n_{\mathrm{E}}$, suppressing power on large scales. For instance, considering the zero conductivity case with $n_{\mathrm{E}}\simeq 0.01$, $T_{\rm re}\simeq 1\,\mathrm{GeV}$, and $w_{re}\simeq 0.5$, the magnetic field today at $1\,\mathrm{Mpc}^{-1}$ is $B_{0}\simeq 3.04\times 10^{-14}\,\mathrm{G}$. In contrast, for the same reheating parameters but in the infinite conductivity case, the present-day magnetic field is significantly weaker, $B_{0}\simeq 3.01\times 10^{-32}\,\mathrm{G}$.
@@
These results highlight the crucial role of post-inflationary gauge field evolution in determining the final amplitude of the magnetic field. In particular, the electrical conductivity of the Universe during reheating has a significant impact on the resulting magnetic field strength observable today.